How do you find the vertex, focus and directrix of a parabola 1/2y^2+18x=0
Answer by lwsshak3(6761)

How do you find the vertex, focus and directrix of a parabola
1/2y^2+18x=0
rewrite equation into standard form: (y-k)^2=-4p(x-h)
y^2/2=-18x
y^2=-36x
This is an equation of a parabola that opens leftwards.
vertex:(0,0)
axis of symmetry: x-axis or y=0
4p=36
p=9
focus(0,-9) (p distance left of vertex on the axis of symmetry)
directrix:(0,9) (p distance right of vertex on the axis of symmetry)